# Determinism and Asymmetry in General Relativity

### v1

@@ 4 De Dicto + Uniqueness
As we have seen, Belot (1995) naturally strengthens de re determinism (his Definition 2) by introducing a uniqueness clause to obtain de re* determinism (his Definition 3). In the analogous way, he could have also considered a natural strengthening of the de dicto form of determinism (his Definition 1). Consider the following.
@@
Definition 5. A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto* deterministic if, for any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then there is a unique isometry $\psi:M\rightarrow M^{\prime}$.
@@
We have seen that within the context of (standard) general relativity, de re determinism and de re* determinism are equivalent. This is not the case for the de dicto and de dicto* types of determinism; we will soon show that the latter condition is strictly stronger than the former. First, we establish that the de re* and de dicto* forms of MLE determinism are independent conditions in the sense that neither implies the other. Consider the following examples.
@@
Example 2. Let $(M,g)$ be four-dimensional Minkowski spacetime. The collection $\mathscr{C}=\{(M,g)\}$ is de re* deterministic as a consequence of Corollary 2. So the collection is de re deterministic from Corollary 1 and hence de dicto deterministic from Proposition 1. But $\mathscr{C}$ cannot be de dicto* deterministic since there are non-trivial isometries $\psi:M\rightarrow M$ of Minkowski spacetime to itself (e.g. rotations, translations).
@@
Example 3. Let $(M,g)$ be two-dimensional Minkowski spacetime in standard $(t,x)$ coordinates. Let $M_{1}$ be the $t<0$ region of $M$ and let $M_{2}$ be the $t<x$ region of $M$. Consider the spacetime $(M^{\prime},g)$ where $M^{\prime}=M_{1}\cup M_{2}$ (see Figure 2). One can verify that $(M^{\prime},g)$ is a globally hyperbolic spacetime and, due to the “missing” region, the only isometry $\psi:M^{\prime}\rightarrow M^{\prime}$ is the identity map. It follows that the collection $\mathscr{C}=\{(M^{\prime},g)\}$ must de dicto* deterministic. But the collection fails to be de re deterministic. To see this, consider the initial segment $U\subset M^{\prime}$ given by the $t<-1$ region. Clearly there is a reflection isometry $\varphi:U\rightarrow U$ defined by $\varphi(t,x)=(t,-x)$. But since the only isometry $\psi:M^{\prime}\rightarrow M^{\prime}$ is the identity map, we find that $\psi_{|U}\neq\varphi$. Since de re determinism is not satisfied, de re* determinism is also not satisfied.
@@
Figure 2: The spacetime $(M^{\prime},g)$ from Example 3. Because of the “missing” region (above the dotted line), the only global isometry is the identity map. But there is a non-trivial reflection isometry from the initial segment $U$ to itself.
@@
Examples 2 and 3 show the unusual character of de dicto* determinism. Example 2 shows a sense in which the definition is quite strong since it will count as indeterministic some collections of spacetimes that might seem intuitively deterministic. Strong as it is, Example 3 shows a sense in which the definition is also quite weak since it does not imply either of the de re forms of MLE determinism. We see that the agreement of the initial segment $U$ with itself given by $\varphi$ does not extend to a global agreement. So the example does not satisfy the de re forms of MLE determinism. Of course, the agreement of the initial segment $U$ with itself given by $\psi_{|U}$ does extend to a global agreement via $\psi$. Moreover, this global agreement is unique which ensures the example satisfies the de dicto* form of MLE determinism. Stepping back, we see that the curious state of affairs obtains because there are non-trivial symmetries of the initial segment $U$ but no such global symmetries. An analogous situation obtains in the collapsing columns example of Wilson (1993). At least, that is the case if the columns collapse in a sufficiently asymmetric way (see Dewar 2024, footnote 28). This suggests that the de dicto* form of MLE is not strong enough by the lights of certain haecceitist positions. In the next section, we will explore the possibility of stronger MLE definitions of determinism that have the effect of ruling out non-trivial local symmetries like $\varphi$. Here, we continue to explore the de dicto* form and its relationship to other determinism definitions.
@@
Recall that Example 1 is the collection $\mathscr{C}=\{(M,g)\}$ where $(M,g)$ is the $t<0$ portion of Minkowski spacetime in standard $(t,x)$ coordinates. Because $\mathscr{C}$ is a singleton collection, it is immediate that it is de dicto deterministic. But it is not de dicto* deterministic since there is a non-trivial global reflection isometry $\psi:M\rightarrow M$ given by $\psi(t,x)=(t,-x)$. We now see that Examples 1, 2, and 3 give us the following.
@@
Proposition 4. If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto* deterministic, then it is de dicto deterministic; the other direction does not hold. Moreover, neither de dicto* determinism nor de re* determinism implies the other condition.
@@
The entire situation so far is captured in the diagram below. Arrows correspond to implication relations. If two conditions in the diagram are not connected by an arrow (or series of arrows), then the corresponding implication relation does not hold. We emphasize that these implication relations hold within (standard) general relativity but not generally (e.g. Leibnizian spacetime is de re but not de re* deterministic).
(de $\begin{array}[]{ccc}\text{(de re*)}&&\text{(de dicto*)}\\ \Updownarrow&&\Downarrow\\ \text{(de re)}&\Rightarrow&\text{(de dicto)}\\ \end{array}$
Why might one be interested in de dicto* determinism? Belot (1995) has argued the de dicto, de re, and de re* forms of MLE determinism all fail to be strong enough to count all instances of indeterminism as such. Near the end of the paper, an example is presented. Belot claims the example is deterministic according to de re* determinism (and hence the de re and de dicto forms as well) and yet it is indeterministic by the lights of certain haecceitist positions (for related examples, see Melia 1999; Pooley 2021). Here is Belot (1995, p. 193):
@@
In this example, $W$ is a world with spacetime points and Newtonian spacetime structure. It initially contains a single $\alpha$ particle. The laws of nature decree that thirteen years later, at $t=0$, the $\alpha$ particle decays into continuum many $\beta$ particles; arranged so that at time $t$, the $\beta$ particles form a spherical shell of radius $t$; with each $\beta$ particle moving away from the center of the sphere along its radius.
@@
A global rotational symmetry of $W$ takes the worldline of one particle ($\beta_{1}$) into the worldline of another ($\beta_{2}$). The presence of such a symmetry supposedly allows one to make sense of the idea that “$\beta_{1}$ could have been $\beta_{2}$” which signals the presence of a type of indeterminism (Belot 1995, p. 194).
One certainly can object to this line of reasoning (Brighouse 1997). Indeed, it has been argued that the example is ambiguous as it stands (Halvorson et al. 2025): one natural way of making it precise renders it de re* deterministic while another way – a less natural approach using names for the $\beta$ particles – renders it de re* indeterministic. We suspect that there will be some philosophers who will nonetheless want to hold that the example is de re* deterministic (as Belot claimed) and yet will still be tempted to argue that “$\beta_{1}$ could have been $\beta_{2}$.” This is nonsense: the example is de re* deterministic only if the $\beta$ particles are not named and “$\beta_{1}$ could have been $\beta_{2}$” makes sense only if $\beta$ particles are named (Halvorson et al. 2025).
@@
Here, we wish to highlight that even if one went along with the nonsensical line that both (i) the example is de re* deterministic and also (ii) “$\beta_{1}$ could have been $\beta_{2}$” is a meaningful statement, then there exist other versions of MLE determinism – e.g. de dicto* – that have the effect of ruling out this type of symmetric example within the context of general relativity. These other definitions are of independent interest since they will allow us to show that general relativity is deterministic in ways that have not been previously appreciated.
@@
We close this section with a question that concerns a connection between de dicto* determinism and the theorem of Choquet-Bruhat and Geroch (1969). Note that because of Example 2 (the singleton collection of Minkowski spacetime), there is no analogue to Corollary 2 for de dicto* determinism: some collections $\mathscr{C}\subseteq\mathscr{V}^{+}$ of four-dimensional, inextendible, globally hyperbolic, vacuum solutions to Einstein’s equation fail to be de dicto* deterministic. Which sub-collections of $\mathscr{V}^{+}$ do count as de dicto* deterministic? Consider the following asymmetry condition (cf. Barrett et al. 2023).
@@
Definition 6. A collection $\mathscr{C}\subseteq\mathscr{U}$ is giraffe if, for any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$, if there are isometries $\varphi,\psi:M\rightarrow M^{\prime}$, then $\varphi=\psi$.
@@
One can verify that any collection $\mathscr{C}\subseteq\mathscr{U}$ that is giraffe must also be rigid. The other direction does not hold since the collection $\mathscr{U}$ is rigid from Proposition 3 but fails to be giraffe since it contains Minkowski spacetime. Moreover, it follows easily that any collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto* deterministic if and only if it is both de dicto deterministic and giraffe. The isometry uniqueness clause which is added to de dicto determinism to obtain de dicto* determinism has the effect of requiring collections of spacetimes be giraffe. In the same way that the rigidity condition is a type of dual to de re determinism, we see that the giraffe condition is a type of dual to de dicto determinism: the latter condition requires the existence of an isometry $\psi:M\rightarrow M^{\prime}$ while the former requires its uniqueness. Because any collection $\mathscr{C}\subseteq\mathscr{H}$ that is both giraffe and de dicto deterministic must be de dicto* deterministic, Proposition 2 implies the following.
Corollary 3. Any giraffe collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de dicto* deterministic.
@@
One naturally wonders: is there a non-empty giraffe collection $\mathscr{C}\subseteq\mathscr{V}^{+}$? Presumably there is. The collection of all spacetimes with trivial global symmetries is a giraffe collection and there is reason to think that its spacetimes are, in some sense, “generic” among the collection $\mathscr{U}$ of all spacetimes (see Mounoud 2015, Theorem 1).
5 De Dicto/De Re + Local Uniqueness
@@
There is a sense in which de dicto* and de re* forms of MLE determinism, even as strong as they are, could still be considered too weak by the lights of certain haecceitist positions. The de dicto* form is considered not strong enough since the classical analogue counts the collapsing columns example as deterministic. That example is counted as indeterministic under the classical analogue of the de re* form. But by the lights of certain (in our view, misguided) haecceitist positions, the de re* definition has its own problems. Belot has argued that it is also not strong enough in the sense that it wrongly counts the particle decay example as deterministic.
@@
In both examples, symmetries are present; local symmetries are present in the initial segment of the collapsing columns example and global symmetries are present in the particle decay example. This naturally suggests a way to strengthen each definition: rule out all symmetries. We shall do this by adding a “local uniqueness” clause to both the de dicto and de re forms of MLE determinism: even parts of worlds (including initial segments) must agree in only one way.
@@
For any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{U}$, let us say that a map $\psi:M\rightarrow M^{\prime}$ is a (not necessarily proper) isometric embedding if $\psi$ is an isometry from $M$ to $\psi[M]$. Consider the following local asymmetry condition (cf. Manchak and Barrett 2023).
@@
Definition 7. A collection $\mathscr{C}\subseteq\mathscr{U}$ is Heraclitus if, for any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any open set $O\subseteq M$, if there are isometric embeddings $\varphi,\psi:O\rightarrow M^{\prime}$, then $\varphi=\psi$.
@@
We see that the Heraclitus condition ensures that even local isometries between spacetimes must be unique. Let $\mathscr{C}\subseteq\mathscr{U}$ be any Heraclitus collection and consider any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$. For any initial segment $U\subset M$, we find there is at most one initial segment $U^{\prime}\subset M^{\prime}$ isometric to $U$. Moreover, if there is an isometry relating $U$ and $U^{\prime}$, it is unique. One can verify that any collection that is Heraclitus must also be giraffe (and thus rigid). The other direction does not hold since the collection $\mathscr{C}=\{(M^{\prime},g)\}$ from Example 3 is giraffe but fails to be Heraclitus. Because the Heraclitus asymmetry condition is the natural strengthening of the giraffe asymmetry condition, we immediately see the natural way to strengthen the de dicto* and de re* forms of determinism – at least within the context of general relativity. We have the following.
Definition 8. A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic if it is de dicto and Heraclitus.
Definition 9. A collection $\mathscr{C}\subseteq\mathscr{H}$ is is de re** deterministic if it is de re and Heraclitus.
@@
Because the Heraclitus condition in each of the definitions ensures that even local isometries between spacetimes must be unique, it serves to rule out even the local version of a “$\beta_{1}$ could have been $\beta_{2}$” argument. Because of Examples 2 and 3 and the way we have set things up, we have the following.
@@
Proposition 5. If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic, then it is de dicto* deterministic (and thus de dicto deterministic); the other direction does not hold. If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de re** deterministic, then it is de re* deterministic (and thus de re deterministic); the other direction does not hold.
@@
Curiously, we find that de dicto** determinism and de re** determinism are equivalent. Consider the following.
Proposition 6. A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic if and only if it is de re** deterministic.
Proof.
@@
One direction is trivial. Suppose a collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic. So $\mathscr{C}$ is Heraclitus. We need only show that it is also de re deterministic. Let $(M,g)$ and $(M^{\prime},g^{\prime})$ be any spacetimes in $\mathscr{C}$ and let $U\subset M$ and $U^{\prime}\subset M^{\prime}$ be any initial segments. Suppose there is an isometry $\varphi:U\rightarrow U^{\prime}$. Since $\mathscr{C}$ is de dicto** deterministic, it is de dicto deterministic. So there is an isometry $\psi:M\rightarrow M^{\prime}$. We find that $\psi_{|U}$ and $\varphi$ are both isometric embeddings of the open set $U$ into $M^{\prime}$. Because $\mathscr{C}$ is Heraclitus, we have $\psi_{|U}=\varphi$. Thus $\mathscr{C}$ is de re deterministic and we are done. ∎
@@
The entire situation is captured in the diagram below. As before, arrows correspond to implication relations. If two conditions in the diagram are not connected by an arrow (or series of arrows), then the corresponding implication relation does not hold. We emphasize that, although these implication relations hold within general relativity, there are open questions here about what the situation looks outside of this context (e.g. various classical spacetime theories).
(de re**)(de $\begin{array}[]{ccc}\text{(de re**)}&\Leftrightarrow&\text{(de dicto**)}\\ \Downarrow&&\Downarrow\\ \text{(de re*)}&&\text{(de dicto*)}\\ \Updownarrow&&\Downarrow\\ \text{(de re)}&\Rightarrow&\text{(de dicto)}\\ \end{array}$
@@
Despite the strength of de dicto**/de re** determinism, these definitions are still weak enough to be satisfied by non-empty collections of spacetimes. Consider the following.
@@
Example 4. Let $(M,g)$ be the two-dimensional spacetime constructed in Manchak and Barrett (2023). This spacetime has the “Heraclitus” property of spacetime which requires that, for any open regions $U,U^{\prime}\subseteq M$, if there is an isometry $\psi:U\rightarrow U^{\prime}$, then $U=U^{\prime}$ and $\psi$ is the identity map. Let $(M^{\prime},g)$ be any globally hyperbolic region of $(M,g)$. (We know that arbitrarily small such regions exist around any point in any spacetime.) It follows easily that $(M^{\prime},g)$ inherits the Heraclitus property of spacetime from $(M,g)$. So $\mathscr{C}=\{(M^{\prime},g)\}$ must be a Heraclitus collection . But $\mathscr{C}$ is also de dicto deterministic. This follows easily since there is only one spacetime in the collection. Since $\mathscr{C}$ is both Heraclitus and de dicto deterministic, it is de dicto** deterministic. Thus, by Proposition 6, it is de re** deterministic as well.

### v2

@@ 4. De Dicto + Uniqueness
As we have seen, Belot ([1995]) naturally strengthens de re determinism (his Definition 2) by introducing a uniqueness clause to obtain de re* determinism (his Definition 3). In the analogous way, he could have also considered a natural strengthening of the de dicto form of determinism (his Definition 1). Consider the following:
@@
Definition 5: A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto* deterministic if, for any spacetimes $(M,g)$, $(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then there is a unique isometry $\psi:M\rightarrow M^{\prime}$.
@@
We have seen that within the context of (standard) general relativity, de re determinism and de re* determinism are equivalent. This is not the case for the de dicto and de dicto* types of determinism; we will soon show that the latter condition is strictly stronger than the former. First, we establish that the de re* and de dicto* forms of MLE determinism are independent conditions in the sense that neither implies the other. Consider the following examples:
@@
Example 2: Let $(M,g)$ be four-dimensional Minkowski spacetime. The collection $\mathscr{C}=\{(M,g)\}$ is de re* deterministic as a consequence of Corollary 2. So the collection is de re deterministic from Corollary 1 and hence de dicto deterministic from Proposition 1. But $\mathscr{C}$ cannot be de dicto* deterministic since there are non-trivial isometries $\psi:M\rightarrow M$ of Minkowski spacetime to itself (e.g. rotations, translations).
@@
Example 3: Let $(M,g)$ be two-dimensional Minkowski spacetime in standard $(t,x)$ coordinates. Let $M_{1}$ be the $t<0$ region of $M$ and let $M_{2}$ be the $t<x$ region of $M$. Consider the spacetime $(M^{\prime},g)$ where $M^{\prime}=M_{1}\cup M_{2}$ (see Figure 2). One can verify that $(M^{\prime},g)$ is a globally hyperbolic spacetime and, due to the ‘missing’ region, the only isometry $\psi:M^{\prime}\rightarrow M^{\prime}$ is the identity map. It follows that the collection $\mathscr{C}=\{(M^{\prime},g)\}$ must be de dicto* deterministic. But the collection fails to be de re deterministic. To see this, consider the initial segment $U\subset M^{\prime}$ given by the t<1t<-1 region. Clearly there is a reflection isometry $\varphi:U\rightarrow U$ defined by $\varphi(t,x)=(t,-x)$. But since the only isometry $\psi:M^{\prime}\rightarrow M^{\prime}$ is the identity map, we find that $\psi_{|U}\neq\varphi$. Since de re determinism is not satisfied, de re* determinism is also not satisfied. (Similar examples showing the satisfaction of de dicto* determinism while violating de re determinism can be constructed even if one were to restrict attention to ‘maximal’ spacetimes. See Manchak ([2016], [2017], [2023]) for more on this spacetime property.)
@@
Figure 2: The spacetime $(M^{\prime},g)$ from Example 3. Because of the ‘missing’ region (above the dotted line), the only global isometry is the identity map. But there is a non-trivial reflection isometry from the initial segment $U$ to itself.
@@
Examples 2 and 3 show the unusual character of de dicto* determinism. Example 2 shows a sense in which the definition is quite strong since it will count as indeterministic some collections of spacetimes that might seem intuitively deterministic. Strong as it is, Example 3 shows a sense in which the definition is also quite weak since it does not imply either of the de re forms of MLE determinism. We see that the agreement of the initial segment $U$ with itself given by $\varphi$ does not extend to a global agreement. So the example does not satisfy the de re forms of MLE determinism. Of course, the agreement of the initial segment $U$ with itself given by $\psi_{|U}$ (the identity map on $U$) does extend to a global agreement via $\psi$ (the identity map on $M^{\prime}$). Moreover, this global agreement is unique which ensures the example satisfies the de dicto* form of MLE determinism.
Stepping back, we see that the curious state of affairs obtains because there are non-trivial symmetries of the initial segment $U$ but no such global symmetries. An analogous situation obtains in the collapsing columns example of Wilson (1993). At least, that is the case if the columns collapse in a sufficiently asymmetric way (see Dewar [2025], footnote 28). This suggests that the de dicto* form of MLE is not strong enough by the lights of certain haecceitist positions. In the next section, we will explore the possibility of stronger MLE definitions of determinism that have the effect of ruling out non-trivial local symmetries like the reflection $\varphi$ in Example 3. Here, we continue to explore the de dicto* form and its relationship to other definitions of determinism.
@@
Recall that Example 1 is the collection $\mathscr{C}=\{(M,g)\}$ where $(M,g)$ is the $t<0$ portion of Minkowski spacetime in standard $(t,x)$ coordinates. Because $\mathscr{C}$ is a singleton collection, it is immediate that it is de dicto deterministic. But it is not de dicto* deterministic since there is a non-trivial global reflection isometry $\psi:M\rightarrow M$ given by $\psi(t,x)=(t,-x)$. We now see that Examples 1, 2, and 3 give us the following:
@@
Proposition 4: If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto* deterministic, then it is de dicto deterministic; the other direction does not hold. Moreover, neither de dicto* determinism nor de re* determinism implies the other condition.
@@
The entire situation so far is captured in the diagram presented here. Arrows correspond to implication relations. If two conditions in the diagram are not connected by an arrow (or series of arrows), then the corresponding implication relation does not hold. We emphasize that these implication relations hold within (standard) general relativity but not generally (e.g. Leibnizian spacetime is de re but not de re* deterministic).
[…]
@@
Why might one be interested in de dicto* determinism? Belot ([1995]) has argued the de dicto, de re, and de re* forms of MLE determinism all fail to be strong enough to count all instances of indeterminism as such. Near the end of the paper, an example is presented. Belot claims the example is deterministic according to de re* determinism (and hence the de re and de dicto forms as well) and yet it is indeterministic by the lights of certain haecceitist positions (for related examples, see Melia [1999]; Pooley [2021]). Here is Belot ([1995], p. 193): ‘In this example, $W$ is a world with spacetime points and Newtonian spacetime structure. It initially contains a single $\alpha$ particle. The laws of nature decree that thirteen years later, at $t=0$, the $\alpha$ particle decays into continuum many $\beta$ particles; arranged so that at time $t$, the $\beta$ particles form a spherical shell of radius $t$; with each $\beta$ particle moving away from the center of the sphere along its radius.’
@@
A global rotational isomorphism of the model $W$ (that preserves both the metric and matter content) takes the worldline of one particle ($\beta_{1}$) into the worldline of another ($\beta_{2}$). The presence of such a symmetry supposedly allows one to make sense of the idea that $\beta_{1}$ could have been $\beta_{2}$’ which signals the presence of a type of indeterminism (Belot [1995], p. 194).
@@
One can object to this line of reasoning (Brighouse [1997]). Indeed, it has been argued that the example is ambiguous as it stands (Halvorson et al. [unpublished]): one natural way of making it precise renders it de re* deterministic while another way – a less natural approach using names for the $\beta$ particles – renders it de re* indeterministic. Here, we wish to highlight that even if one holds that both (i) the example is de re* deterministic and also (ii) $\beta_{1}$ could have been $\beta_{2}$’ is a meaningful statement, then there exist other versions of MLE determinism – e.g. de dicto* – that have the effect of ruling out this type of symmetric example within the context of general relativity. These other definitions are of independent interest since they will allow us to show that general relativity is deterministic in ways that have not been previously appreciated.
@@
We close this section with a question that concerns a connection between de dicto* determinism and the theorem of Choquet-Bruhat and Geroch ([1969]). Note that because of Example 2 (the singleton collection of Minkowski spacetime), there is no analogue to Corollary 2 for de dicto* determinism: some collections $\mathscr{C}\subseteq\mathscr{V}^{+}$ of four-dimensional, inextendible, globally hyperbolic, vacuum solutions to Einstein’s equation fail to be de dicto* deterministic. So which sub-collections of $\mathscr{V}^{+}$ do count as de dicto* deterministic? Consider the following asymmetry condition (Barrett et al. [2023]):
@@
Definition 6: A collection $\mathscr{C}\subseteq\mathscr{U}$ is giraffe if, for any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$, if there are isometries $\varphi,\psi:M\rightarrow M^{\prime}$, then $\varphi=\psi$.
@@
One can verify that any collection $\mathscr{C}\subseteq\mathscr{U}$ that is giraffe must also be rigid. The other direction does not hold since the collection $\mathscr{U}$ is rigid from Proposition 3 but fails to be giraffe since it contains Minkowski spacetime. Moreover, it follows easily that any collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto* deterministic if and only if it is both de dicto deterministic and giraffe. To sum up: the isometry uniqueness clause which is added to de dicto determinism to obtain de dicto* determinism has the effect of requiring collections of spacetimes to be giraffe.
In the same way that the rigidity condition is a type of dual to de re determinism, we see that the giraffe condition is a type of dual to de dicto determinism: the latter condition requires the existence of an isometry $\psi:M\rightarrow M^{\prime}$ while the former requires its uniqueness. Because any collection $\mathscr{C}\subseteq\mathscr{H}$ that is both giraffe and de dicto deterministic must be de dicto* deterministic, Proposition 2 implies the following:
Corollary 3: Any giraffe collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de dicto* deterministic.
@@
One naturally wonders: is there a non-empty giraffe collection $\mathscr{C}\subseteq\mathscr{V}^{+}$? Presumably there is although we are not aware of any result showing this. It is easy to come up with an example that is a four-dimensional, globally hyperbolic, vacuum solution with trivial global symmetries by removing certain regions from Minkowski spacetime. But such an example fails to be inextendible. However, the collection of all spacetimes with trivial global symmetries (i.e. only the identity map as a symmetry) is a giraffe collection and there is reason to think that its spacetimes are, in some sense, ‘generic’ among the collection $\mathscr{U}$ of all spacetimes (see Mounoud [2015], Theorem 1). Perhaps there is a sense in which such spacetimes are generic among the collection $\mathscr{V}^{+}$ as well.
5. De Dicto/De Re + Local Uniqueness
@@
There is a sense in which de dicto* and de re* forms of MLE determinism, strong though they are, could still be considered too weak by the lights of certain haecceitist positions. The de dicto* form is considered not strong enough since the classical analogue counts the collapsing columns example as deterministic. That example is counted as indeterministic under the classical analogue of the de re* form. But by the lights of certain (in our view, misguided) haecceitist positions, the de re* definition has its own problems. Belot has argued that it is also not strong enough in the sense that it wrongly counts the particle decay example as deterministic.
@@
In both examples, symmetries are present; local symmetries are present in the initial segment of the collapsing columns example and global symmetries are present in the particle decay example. This naturally suggests a way to strengthen each definition: rule out all symmetries. We shall do this by adding a ‘local uniqueness’ clause to both the de dicto and de re forms of MLE determinism: even parts of worlds (including initial segments) must agree in only one way.
@@
For any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{U}$, let us say that a map $\psi:M\rightarrow M^{\prime}$ is a (not necessarily proper) ‘isometric embedding’ if $\psi$ is an isometry from $M$ to $\psi[M]$. Consider the following local asymmetry condition (Manchak and Barrett [forthcoming]):
@@
Definition 7: A collection $\mathscr{C}\subseteq\mathscr{U}$ is Heraclitus if, for any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any open set $O\subseteq M$, if there are isometric embeddings $\varphi,\psi:O\rightarrow M^{\prime}$, then $\varphi=\psi$.
@@
The Heraclitus condition ensures that even local isometries between spacetimes must be unique. Let $\mathscr{C}\subseteq\mathscr{U}$ be any Heraclitus collection and consider any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$. For any initial segment $U\subset M$, we find there is at most one initial segment $U^{\prime}\subset M^{\prime}$ isometric to $U$. Moreover, if there is an isometry relating $U$ and $U^{\prime}$, it is unique. One can verify that any collection that is Heraclitus must also be giraffe (and thus rigid). The other direction does not hold since the collection $\mathscr{C}=\{(M^{\prime},g)\}$ from Example 3 is giraffe but fails to be Heraclitus. Because the Heraclitus asymmetry condition is the natural local strengthening of the giraffe asymmetry condition, we immediately see the natural way to strengthen the de dicto* and de re* forms of determinism – at least within the context of general relativity. We have the following:
Definition 8: A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic if it is de dicto and Heraclitus.
Definition 9: A collection $\mathscr{C}\subseteq\mathscr{H}$ is is de re** deterministic if it is de re and Heraclitus.
@@
Because the Heraclitus condition in each of the definitions ensures that even local isometries between spacetimes must be unique, it serves to rule out even the local version of a $\beta_{1}$ could have been $\beta_{2}$’ argument. Because of Examples 2 and 3 and the way we have set things up, we have the following:
@@
Proposition 5: If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic, then it is de dicto* deterministic (and thus de dicto deterministic); the other direction does not hold. If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de re** deterministic, then it is de re* deterministic (and thus de re deterministic); the other direction does not hold.
@@
Curiously, we find that de dicto** determinism and de re** determinism are equivalent. Consider the following:
Proposition 6: A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic if and only if it is de re** deterministic.
@@
Proof: One direction is trivial. Suppose a collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto** deterministic. So $\mathscr{C}$ is Heraclitus. We need only show that it is also de re deterministic. Let $(M,g)$ and $(M^{\prime},g^{\prime})$ be any spacetimes in $\mathscr{C}$ and let $U\subset M$ and $U^{\prime}\subset M^{\prime}$ be any initial segments. Suppose there is an isometry $\varphi:U\rightarrow U^{\prime}$. Since $\mathscr{C}$ is de dicto** deterministic, it is de dicto deterministic. So there is an isometry $\psi:M\rightarrow M^{\prime}$. We find that $\psi_{|U}$ and $\varphi$ are both isometric embeddings of the open set $U$ into $M^{\prime}$. Because $\mathscr{C}$ is Heraclitus, we have $\psi_{|U}=\varphi$. Thus $\mathscr{C}$ is de re deterministic and we are done. ∎
@@
The entire situation is captured in the diagram presented here. As before, arrows correspond to implication relations. If two conditions in the diagram are not connected by an arrow (or series of arrows), then the corresponding implication relation does not hold. We emphasize that, although these implication relations hold within general relativity, there are open questions here about what the situation looks outside of this context. (All down arrows and right arrows depicted in the diagram hold in general. But as we have seen, Leibnizian spacetime satisfies the classical analogue to de re determinism but not de re* determinism. One wonders about the situation regarding de dicto** and de re** determinism within the context of various classical spacetime theories. Do classical Heraclitus spacetimes even exist?)
[…]
@@
Despite the strength of de dicto**/de re** determinism, these definitions are still weak enough to be satisfied by non-empty collections of spacetimes. Consider the following:
@@
Example 4: Let $(M,g)$ be the two-dimensional spacetime constructed in Manchak and Barrett ([forthcoming], pp. 11-14). This spacetime has the ‘Heraclitus’ property of spacetime which requires that, for any open regions $U,U^{\prime}\subseteq M$, if there is an isometry $\psi:U\rightarrow U^{\prime}$, then $U=U^{\prime}$ and $\psi$ is the identity map. The example is conformally flat with a manifold diffeomorphic to $\mathbb{R}^{2}$. The Heraclitus property obtains because of a careful choice of conformal factor which ensures that no distinct points in $M$ have all of the same scalar curvature values. Let $(M^{\prime},g)$ be any globally hyperbolic region of $(M,g)$. (We know that arbitrarily small such regions exist around any point in any spacetime.) It follows easily that $(M^{\prime},g)$ inherits the Heraclitus property of spacetime from $(M,g)$. So $\mathscr{C}=\{(M^{\prime},g)\}$ must be a Heraclitus collection. But $\mathscr{C}$ is also de dicto deterministic. This follows easily since there is only one spacetime in the collection. Since $\mathscr{C}$ is both Heraclitus and de dicto deterministic, it is de dicto** deterministic. Thus, by Proposition 6, it is de re** deterministic as well.